26. Calculate the resultant of the four forces working on an object:

25 N 20° W of N

60 N 18° S of W

45 N W

35 N 49° E of N

Use the perpendicular component method. Express the answer in unit vector notation.  (6)



1
Expert's answer
2021-06-13T07:09:43-0400

"F_x=-25\cdot\sin20°-45\cdot\cos45°-60\cdot\cos18°+35\cdot\sin49°\approx-71\ (N)"


"F_y=25\cdot\cos20°+45\cdot\sin45°-60\cdot\sin18°+35\cdot\cos49°\approx60\ (N)"


"F=\sqrt{71^2+60^2}\approx93\ (N)"


"\vec F=-71\vec i+60\vec j" . Answer